On Monday, Roger drove to work in 55 minutes. On Tuesday, he averaged 8 miles per hour more, and it took him 8 minutes less to get to work. How far (in miles) does he travel to work? : On Monday, Roger drove to work in 55 minutes. On Tuesday, he averaged 8 miles per hour more, and it took him 8 minutes less to get to work. How far (in miles) does he travel to work?
Answer by ankor@dixie-net.com(3949)

On Monday, Roger drove to work in 55 minutes. On Tuesday, he averaged 8 miles per hour more, and it took him 8 minutes less to get to work. How far (in miles) does he travel to work?
:
Change 55 min to hrs, 55/60 hrs, time on Monday
Change 8 min to hrs, 8/60
55/60 - 8/60 = 47/60hrs, time on Tuesday
:
First, find the speed:
s = speed on Monday
then
(s+8) = speed on Tuesday
:
The distance of the two trips is equal;
write a distance equation: Dist = time * speed.
:
55/60s = (47)/60(s+8)
:
Multiply both sides by 60, get rid of the denominators:
55s = 47(s+8)
55s = 47s + 376
55s - 47s = 376
8s = 376
s = 376/8
s = 47 mph, speed on Monday
then
47 + 8 = 55 mph speed on Tuesday
:
Dist using Monday's time & speed
Dist = 55/60 * 47
Dist = 43.1 mi
:
Check dist using Tues time & speed
Dist = 47/60 * (47+8)
Dist = 47/60 * 55
Dist = 43.1 mi